# On relative tractor bundles

### v1

@@ 3.5. Example: Generalized path geometries
(1) Similarly as in the case of Legendrean contact structures, Theorem 12 covers many cases in which no good description of the sheaves resolved by relative BGG sequences was known. Indeed, such a description was only known in the case of a correspondence space treated in part (2) of Theorem 4.11 of [Rel-BGG2]. In the case of generalized path geometries (with involutive subbundle $V$) these are exactly those geometries for which the harmonic curvature component of homogeneity $3$ mentioned in the proof of Theorem 12 vanishes identically. This means that locally the family of paths can be realized as geodesics of a torsion-free linear connection on the tangent bundle of a local leaf space. Since the unparametrized geodesics depend only on the projective equivalence class of the connection, such correspondence spaces in dimension $12n+$ 1 are best described as induced from a projective structure in dimension $n+1$. Of course, there are many systems of second order ODE which do not admit a description via geodesics, and these lead to generalized path geometries for which the harmonic curvature component of homogeneity $3$ is non-trivial.
@@
(2) As remarked already, the situation is not as symmetric here as for Legendrean contact structures. However, from the point of view of this article, the pair $\mathfrak{q}\subset\tilde{\mathfrak{p}}\subset\mathfrak{g}$ is much less interesting. This leads to $T_{\rho}M=E$ and the basic relative tractor bundle $TM/E$. On the one hand, the fact that $T_{\rho}M$ has rank one, then implies that all relative BGG sequences consist of a single operator only. On the other hand, the class of relative tractor bundles provided by our construction is less rich then in the cases discussed in Theorems 10 and 12. The point here is that essentially $\tilde{P}/\tilde{P}_{+}\cong S(GL(2,\mathbb{R})\times GL(n,\mathbb{R}))\subset GL(2n,\mathbb{R})$ in our case and the bundle $TM/E$ corresponds to the tensor product of the dual of the standard representation of the $GL(2,\mathbb{R})$ factor with the standard representation of the $GL(n,\mathbb{R})$-factor. The latter give rise to relative tractor bundles of rank $2$ respectively $n$, that cannot be obtained by tensorial constructions form $TM/E$. The tensor product decomposition does not descend to a local leaf space for $E$ unless the initial geometry is torsion-free, in which case it induces a $(2,n)$-Grassmannian structure on the local leaf space.
@@
(3) The relation of relative tractor bundles to $TM/V$ in the case of generalized path geometries was observed during the work on the thesis [Zhangwen-thesis] of the second author under the direction of the first author. Relative tractor bundles and the relative BGG machinery play only a minor role in the thesis, however. One of the results of the thesis is a complete description of Weyl structures for generalized path geometries and relative tractor connections play an important role in the characterization of Weyl connections and in the description of tractor calculus in terms of Weyl connections. In particular, this also leads to a description of relative tractor bundles and relative tractor connections as discussed in Remark 6.

### v2

@@ 3.5 Example: Generalized path geometries
    Similarly as in the case of Legendrean contact structures, Theorem 3.8 covers many cases in which no good description of the sheaves resolved by relative BGG sequences was known. Indeed, such a description was only known in the case of a correspondence space treated in part (2) of [10, Theorem 4.11]. In the case of generalized path geometries (with involutive subbundle $V$) these are exactly those geometries for which the harmonic curvature component of homogeneity $3$ mentioned in the proof of Theorem 3.8 vanishes identically. This means that locally the family of paths can be realized as geodesics of a torsion-free linear connection on the tangent bundle of a local leaf space. Since the unparametrized geodesics depend only on the projective equivalence class of the connection, such correspondence spaces in dimension $12n+$ 1 are best described as induced from a projective structure in dimension $n+1$. Of course, there are many systems of second order ODE which do not admit a description via geodesics, and these lead to generalized path geometries for which the harmonic curvature component of homogeneity $3$ is non-trivial.
  2. (2)
@@
    As remarked already, the situation is not as symmetric here as for Legendrean contact structures. However, from the point of view of this article, the pair $\mathfrak{q}\subset\tilde{\mathfrak{p}}\subset\mathfrak{g}$ is much less interesting. This leads to $T_{\rho}M=E$ and the basic relative tractor bundle $TM/E$. On the one hand, the fact that $T_{\rho}M$ has rank one, then implies that all relative BGG sequences consist of a single operator only. On the other hand, the class of relative tractor bundles provided by our construction is less rich then in the cases discussed in Theorems 3.6 and 3.8. The point here is that essentially $\tilde{P}/\tilde{P}_{+}\cong S({\rm GL}(2,\mathbb{R})\times{\rm GL}(n,\mathbb{R}))\subset{\rm GL}(2n,\mathbb{R})$ in our case and the bundle $TM/E$ corresponds to the tensor product of the dual of the standard representation of the ${\rm GL}(2,\mathbb{R})$ factor with the standard representation of the ${\rm GL}(n,\mathbb{R})$-factor. The latter give rise to relative tractor bundles of rank $2$ respectively $n$, that cannot be obtained by tensorial constructions form $TM/E$. The tensor product decomposition does not descend to a local leaf space for $E$ unless the initial geometry is torsion-free, in which case it induces a $(2,n)$-Grassmannian structure on the local leaf space.
  3. (3)
@@
    The relation of relative tractor bundles to $TM/V$ in the case of generalized path geometries was observed during the work on the thesis [15] of the second author under the direction of the first author. Relative tractor bundles and the relative BGG machinery play only a minor role in the thesis, however. One of the results of the thesis is a complete description of Weyl structures for generalized path geometries and relative tractor connections play an important role in the characterization of Weyl connections and in the description of tractor calculus in terms of Weyl connections. In particular, this also leads to a description of relative tractor bundles and relative tractor connections as discussed in Remark 3.2.
3.6 Explicit examples of relative BGG resolutions
Here we give a few explicit examples of relative BGG resolutions for the examples treated in the last two sections. Both cases are closely related to standard BGG sequences associated to projective structures in appropriate dimension. We do not go into the details of how to compute the relevant relative Lie algebra homology groups but refer to [9] and [10]. We focus on the case of generalized path geometries in dimension $7$. This is quite close to the five-dimensional case discussed in [10] but shows a bit more varied behavior in simple cases. In the end, we briefly comment on the case of Legendrean contact structures.
[…]
The full relative BGG sequence for $k=1$, i.e., the bundle $(TM/V)^{*}$, has the form
[…]
Up to a twist by a line bundle, the last two bundles can be interpreted as the kernel of the complete symmetrization in $V^{*}\otimes S^{2}V^{*}$ and as $V^{*}$, respectively.
[…]
Up to a twist by a density bundle, the first two bundles in the sequence are induced by the representations $V^{*}$ and $S^{2}V^{*}$, so the first operator is like a symmetrized covariant derivative in $V$-directions. By Theorem 3.8, locally its kernel is isomorphic to the pullbacks of $2$-forms on a local leaf space. Up to a twist by a density bundle, the third bundle is induced by $S^{2}V$, and the second operator is of order two. This operator is the relative analog of the projectively invariant version of the operator that computes the infinitesimal change of curvature caused by an infinitesimal deformation of a Riemannian metric. The last bundle is (up to a twist) induced by $V$ and the last operator roughly is a divergence in $V$-directions. To make the operators in all the BGG sequences more explicit, one has to use Weyl structures for path geometries, see [15].
The relative BGG sequences for Legendrean contact structures as discussed in Section 3.4 are also similar in form to standard BGG sequences for projective structures. If the original manifold $M$ has dimension $12n+$ 1, then relative BGG sequences are related to BGG sequences for projective structures in dimension $n$. Here the basic relative tractor bundle $TM/F$ contains the distinguished subbundle $H/F\cong E\subset TM/F$ of rank $n$, and is corresponds to the dual of the standard tractor bundle in projective geometry. So in particular, the relative BGG sequence determined by this bundle starts with a second order operator defined on a line bundle, and then extends similar to (3.3) (up to twists by appropriate line bundles) with $S^{2}E^{*}$, the kernel of the complete symmetrization in $E^{*}\otimes S^{2}E^{*}$, and $E^{*}$. The operators in the sequence have orders $2$, $1$, and $1$. Again, they can be made more explicit in terms of (families of) distinguished connections, see [22].
Acknowledgements
For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. The first and second authors supported by the Austrian Science Fund (FWF), doi:10.55776/P33559. The second and third authors partially supported by the Austrian Science Fund (FWF), doi:10.55776/Y963. This article is based upon work from COST Action CaLISTA CA21109 supported by COST (European Cooperation in Science and Technology, https://www.cost.eu). We thank the anonymous referees for several helpful comments and suggestions.